object traveling at 90% speed of light what would it appear as to observer from sending point ,the object ,and destination point, and from a 4th observation from 90% angle, 4th point is most important and observed effects of movement on spacetime from 4th point...

1
Expert's answer
2022-04-14T16:01:42-0400

"l=l_0\sqrt{1-\frac{v^2}{c^2}}=0.44l_0."